OpenStudy (anonymous):

The region bounded by the parabola y=x^2 and above by the line y=x-2 Density = 8x^2. Find the center of mass of the thin plate covering the given region

OpenStudy (amistre64):

any ideas?

OpenStudy (anonymous):

LOL what?

OpenStudy (amistre64):

you asked a question, im asking what ideas you have come up with for a solution process.

OpenStudy (anonymous):

OOO ok

OpenStudy (anonymous):

I have that M= \[\int\limits_{-1}^{2} (8x^2)[(x+2)^2-(x^2)^2]\]

OpenStudy (amistre64):

hmm, i was considering a multiple integation setup, something like:\[\int_D\rho(x,y)~dA\]

OpenStudy (amistre64):

limits of y = x^2 to x-2 x = intersectiions

OpenStudy (amistre64):

does that sound like the same process?

OpenStudy (amistre64):

youre curves dont intersect, are they written correctly?

OpenStudy (anonymous):

o its y=x+2 sorry

OpenStudy (amistre64):

\[\int \limits^{x=2}_{x=-1}~\int\limits^{y=x+2}_{y=x^2}8x^2~dy~dx\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I havent gotten to double intergals yet

OpenStudy (amistre64):

well, its just working inside out 8(x^3)/3 from x^2 to x+2 8((x+2)^3)/3 - 8(x^6)/3 that takes care of dy, then its just the usual dx stuff

OpenStudy (amistre64):

pfft, been awhile, pretty sure i messed up

OpenStudy (amistre64):

there are no ys to integrate, so its just y'=1 which integrates to y so 8x^2[(x+2) - (x^2)] dx

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

which is just a product of density times the difference in our functions

OpenStudy (anonymous):

OOOOOOOO

OpenStudy (amistre64):

so out total mass seems to be \[8\int_{-1}^{2}x^3+2x^2-x^4dx=63/20\]for the total mass

OpenStudy (amistre64):

126/5 i forgot the 8 out front lol

OpenStudy (anonymous):

ooooooooo ok

OpenStudy (anonymous):

thats where I was going with my intergration

OpenStudy (amistre64):

i was wondering why you had squared your functions is all

OpenStudy (amistre64):

do you agree that the center of mass will balance out to an average x and y point? if so, then its really equating: \[8\int_{-1}^{b}x^3+2x^2-x^4dx=\frac{126}{2(5)}\]

OpenStudy (amistre64):

where b is the balance point for x

OpenStudy (anonymous):

O ok

OpenStudy (amistre64):

-b^5/5+b^4/4+2b^3/3 - (-(-1)^5/5+(-1)^4/4+2(-1)^3/3) = 126/80 -b^5/5+b^4/4+2b^3/3 - (1/5+1/4-2/3) = 126/80 which works out to about 1.27118 if my idea is good, unless you can think of a formula for it

OpenStudy (anonymous):

thanks so much!

OpenStudy (amistre64):

youre welcome, make sure you chk with your course material to be sure that my idea is sound. all i got to go on is memory at the moment

OpenStudy (amistre64):

finding the average y part of the point would be something similar but trickier since my idea would have to try to find inverses to play with

OpenStudy (amistre64):